# Reliable Grid Forecasting: State Space Models for Safety-Critical Energy Systems

### v1

@@ Abstract
Using this framework, we conduct a systematic evaluation of Mamba-based State Space Models for California grid forecasting on a weather-aligned CAISO TAC-area dataset spanning Nov 2023–Nov 2025 (84,498 hourly records across 5 transmission areas). Our analysis reveals that standard accuracy metrics are poor proxies for operational safety: models with identical MAPE can require vastly different reserve margins.
We demonstrate that forecast errors are weakly but significantly associated with temperature (, ), motivating weather-aware modeling rather than loss function modification alone. The S-Mamba model achieves the lowest Reserve margin (14.12%) compared to 16.66% for iTransformer, demonstrating superior forecast reliability under a 99.5th-percentile tail-risk reserve proxy.
@@
Contributions.
The main contributions of this work are summarized as follows:
@@
Loss vs. Safety Analysis. We provide evidence that loss function modification cannot substitute for proper covariate modeling—quantile loss training reduces large errors by only 1.6% on CA ISO-TAC test windows, while weather effects depend on architecture and matched evaluation alignment.
@@ 6.3 Error Analysis: The Impact of Weather on Forecast Reliability
Temperature-Error Association. Using all horizons of the CA ISO-TAC sliding-window forecasts (113,376 hourly prediction points), we find a statistically significant but modest association between temperature and absolute forecast error (Pearson , ). The fitted slope is 24.1 MW/°C (Fig. 2a). Hot conditions (°C) exhibit a modest increase in mean error (+3.4%).
@@ 7.1 Performance Comparison with CAISO Benchmarks
The performance gains are particularly significant given the model efficiency: PowerMamba achieves CAISO-beating accuracy with only 2.5M parameters—69% fewer than Chronos and 95% fewer than the iTransformer baseline used in our experiments. This efficiency enables real-time deployment on edge devices without sacrificing forecast quality.
@@ 7.2 Temperature and Forecast Errors
Our error analysis reveals a statistically significant but modest association between forecast errors and temperature extremes (, ). Hot conditions above 30°C are associated with a modest increase in mean error. While this association is not large, it supports the operational intuition that weather is a relevant covariate, and that loss function modifications cannot substitute for missing covariate information.
@@ 8 Conclusion
Mamba-based state space models achieve competitive accuracy for California grid load forecasting. With weather integration, PowerMamba achieves 4.11% MAPE for 24-hour forecasts—a 9.6% improvement over CAISO’s operational benchmark (4.55% MAPE). Critically, S-Mamba achieves the lowest Reserve margin (14.12%) among all models, 43.7% lower than the LSTM baseline (25.07%), demonstrating that careful model design can reduce tail-risk reserve requirements while maintaining forecast quality.

### v2

@@ Abstract
Using this framework, we conduct a systematic evaluation of Mamba-based State Space Models for California grid forecasting on a weather-aligned CA ISO-TAC dataset spanning Nov 2023–Nov 2025 (84,498 hourly records across 5 transmission areas). Our analysis reveals that standard accuracy metrics are poor proxies for operational safety: models with identical MAPE can require vastly different reserve margins.
We demonstrate that forecast errors are weakly but statistically significantly associated with temperature (), motivating weather-aware modeling rather than loss function modification alone. The S-Mamba model achieves the lowest Reserve margin (14.12%) compared to 16.66% for iTransformer, demonstrating superior forecast reliability under a 99.5th-percentile tail-risk reserve proxy.
@@
Contributions
To move beyond symmetric accuracy metrics toward operator-relevant reliability for California grid forecasting, we make the following contributions:
@@
Loss vs. Safety Analysis. We show that loss reweighting alone cannot substitute for covariate modeling: quantile loss reduces 1000 MW errors by only 1.6% on CA ISO-TAC test windows, while weather-aware modeling yields materially larger tail-risk improvements (Section 6).
@@ 6.3 Error Analysis: The Impact of Weather on Forecast Reliability
Temperature-Error Association. Using all horizons of the CA ISO-TAC sliding-window forecasts (113,376 hourly prediction points), we find a statistically significant but modest association between temperature and absolute forecast error (Pearson ). The fitted slope is 24.1 MW/°C (Fig. 2a). Hot conditions (°C) exhibit a modest increase in mean error (+3.4%).
@@ 6.5 Weather Integration Results
6.6 NEMs (BtM Registry) Integration Results
We next evaluate whether static NEMs/registry-derived BtM features (e.g., installed PV and storage capacity) provide incremental predictive value on top of the weather-aligned baseline. This analysis is reported only for utilities where NEMs/registry features were available in our pipeline (PGE, SCE, SDGE, TIDC).
How NEMs features are integrated. NEMs/registry features are static or slow-moving, so we integrate them as an additional BtM input stream that is fused into the weather-integrated forecaster. To avoid physically irrelevant feature injection (e.g., PV capacity “explaining” nighttime net load), we apply a simple daylight prior (implemented as a time-of-day gate). When weather inputs are present, we further apply a lightweight context-conditioned modulation so the BtM contribution can vary smoothly by regime (e.g., daylight/irradiance conditions) rather than acting as an unconditional bias.
Concretely, let denote the main embedded inputs (net load history with weather and time features), and let denote the NEMs/registry feature vector. We compute a BtM embedding and modulate it using a context signal (weather+time) via FiLM-style parameters:
then fuse additively before the sequence encoder:
Figure 6 summarizes this “weather + NEMs” fusion path.
Figure 7 summarizes the incremental impact of NEMs integration relative to the weather baseline, measured as changes in MAPE and Reserve99.5 (tail under-forecast requirement proxy). Improvements are heterogeneous: in the PV-dominant reference utility with NEMs availability (PGE), NEMs integration reduces both average error and tail-risk reserve requirement, while other territories show smaller changes and one (SCE) degrades slightly. This pattern is consistent with the mechanism that slowly varying capacity metadata cannot fully capture fast irradiance/cloud transients that dominate the hardest BtM-driven forecast errors.
@@ 7.1 Performance Comparison with CAISO Benchmarks
The performance gains are particularly significant given the model efficiency: PowerMamba achieves competitive accuracy relative to published CAISO benchmarks with only 2.5M parameters—69% fewer than Chronos and 95% fewer than the iTransformer baseline used in our experiments. This efficiency enables real-time deployment on edge devices without sacrificing forecast quality.
@@ 7.2 Temperature and Forecast Errors
Our error analysis reveals a statistically significant but modest association between forecast errors and temperature extremes (). Hot conditions above 30°C are associated with a modest increase in mean error. While this association is not large, it supports the operational intuition that weather is a relevant covariate, and that loss function modifications cannot substitute for missing covariate information.
@@ 7.5 Limitations
NEMs/BtM feature limitations. Static NEMs/registry capacity features provide only modest incremental improvements beyond weather integration and can be heterogeneous across utilities (Section 6.6). This suggests that richer BtM drivers (irradiance alignment, orientation distributions, and/or satellite nowcasting) and auditable gross/solar/storage decomposition are likely necessary for larger, reliable gains.
@@ 8 Conclusion
Mamba-based state space models achieve competitive accuracy for California grid load forecasting. With weather integration, PowerMamba achieves 4.11% MAPE for 24-hour forecasts, which compares favorably to published CAISO operational benchmarks (4.55% MAPE; contextual reference). Critically, S-Mamba achieves the lowest Reserve margin (14.12%) among all models, 43.7% lower than the LSTM baseline (25.07%), demonstrating that careful model design can reduce tail-risk reserve requirements while maintaining forecast quality.